# WiFi-Diffusion: Achieving Fine-Grained WiFi Radio Map Estimation with Ultra-Low Sampling Rate by Diffusion Models

### v1

@@ VI Conclusion
Fine-grained radio maps present communication parameters of interest, e.g., received signal strength, across a large geographical region in a specific frequency band. It can be leveraged to improve the efficiency of spectrum utilization, particularly critical for the WiFi spectrum used by numerous devices for information transmission. The problem of fine-grained radio map estimation involves utilizing samples from sparsely distributed sensors to infer the parameter of interest at every point in the large region. This problem is uniquely challenging due to the ultra-low sampling rate, where the number of radio samples is significantly smaller than the fine-grained resolution required for radio map estimation.
@@
In this paper, we design WiFi-Diffusion – a novel generative framework for fine-grained WiFi radio map estimation using diffusion models. WiFi-Diffusion consists of three important blocks: 1) a boost block, which uses prior information such as the layout of obstacles and laws of radio propagation to optimize the diffusion model; 2) a generation block, which employs the diffusion model to generate a candidate set of radio maps; 3) an election block, which utilizes the mathematical radio propagation model as a guide to find the best radio map from the candidate set. Extensive simulations demonstrate that 1) WiFi-Diffusion has the capability of constructing high-quality, fine-grained radio maps when the sampling rate is ultra-low (less than $0.1\%$); 2) the radio map generated by WiFi-Diffusion can be an order of magnitude (10 times) better than that generated by SOTA when they share the same sampling rate; 3) WiFi-Diffusion requires only one-fifth of the sampling rate needed by SOTA to generate comparable radio maps.

### v3

@@ VII Conclusion
Radio maps present communication parameters of interest, e.g., received signal strength, across a geographical region in a specific frequency band. It can be used to improve the efficiency of spectrum utilization, particularly critical for the WiFi spectrum. The problem of fine-grained radio map estimation involves utilizing samples from sparsely distributed sensors to infer a high-resolution radio map. This problem is uniquely challenging due to the ultra-low sampling rate, where the number of radio samples is significantly smaller than the high resolution required for radio map estimation.
@@
In this paper, we design WiFi-Diffusion – a novel generative framework for fine-grained WiFi radio map estimation using diffusion models. WiFi-Diffusion consists of three blocks: 1) a boost block, which uses prior information such as the layout of obstacles and laws of radio propagation to optimize the diffusion model; 2) a generation block, which employs the diffusion model to generate a candidate set of fine-grained radio maps; 3) an election block, which utilizes the mathematical radio propagation model as a guide to find the best fine-grained radio map from the candidate set. Extensive simulations show that 1) WiFi-Diffusion constructs high-quality, fine-grained radio maps when the sampling rate is extremely low (less than $0.1\%$); 2) the radio map generated by WiFi-Diffusion can be an order of magnitude (10 times) better than that generated by SOTA when they share the same sampling rate; 3) WiFi-Diffusion requires only one-fifth of the sampling rate needed by SOTA to generate comparable radio maps.